there are three consecutive odd integers and sum of smaller is 49 less than 3 times largest integer find the three odd integer
Answer by oberobic(2304)

3 consecutive odd integers: x, x+2, x+4.
.
x + x+2 = 3(x+4) -49
.
2x +2 = 3x +12 -49
-x = -39
x = 39
.
Always check your work.
.
x + x+2 = 39 + 39 + 2
= 80
.
3*(x+4) = 3(43)
= 129
- 49
= 80
Correct.
.
Answer:
The three consecutive odd integers are: 39, 41, and 43.
.
Done.